OpenStudy (pokemon23):

Approximate ,correct to the nearest truth, the third side of the right triangle whose hypotenuse is represented by c and whose other sides are represented by a and b.

OpenStudy (pokemon23):

c=35,b=25

OpenStudy (akshay_budhkar):

is this correct?

OpenStudy (akshay_budhkar):

btw hello pokemon ready for some work?

OpenStudy (pokemon23):

sure help with my math hw

OpenStudy (akshay_budhkar):

ok i am there for good time today. lets begin

OpenStudy (pokemon23):

ok

OpenStudy (akshay_budhkar):

Pythagoras theorem is \[c^2 = a^2 + b^2\]

OpenStudy (akshay_budhkar):

so just plug in the values in the formula above and solve and show me

OpenStudy (pokemon23):

a^2+25^2=35^2

OpenStudy (akshay_budhkar):

go ahead

OpenStudy (pokemon23):

a^2+625=1225

OpenStudy (akshay_budhkar):

ok correct

OpenStudy (pokemon23):

a^2=600

OpenStudy (akshay_budhkar):

yes!

OpenStudy (pokemon23):

24.49489743

OpenStudy (akshay_budhkar):

yea correct :D

OpenStudy (pokemon23):

a=24.49489743

OpenStudy (pokemon23):

nearest tenth

OpenStudy (pokemon23):

how do i get this to the nearest tenth

OpenStudy (akshay_budhkar):

ok nearest tenth is 24.5

OpenStudy (akshay_budhkar):

you round it to the closest tenth, so 24.49.... is closer to 24.5 than it is to 24.4

OpenStudy (pokemon23):

ok second question Find x and express irrational results in simplest radical form.

OpenStudy (pokemon23):